Quali-Rags has fixed costs of $300,000, sells its units for $75, and has variable costs of $50 per unit. Compute the break-even point in units. a. An analyst comes up with a new plan to cut fixed costs to $250,000. However, more raw material will now be required, which will increase variable costs per unit to $59. The sales price will remain at $75. What is the new break-even point in units b. c. Quali-Rags management likes the new plan, but also wants to generate an additional $150,000 in gross profit to purchase some new equipment. Calculate the new break-even point in units using the new plan with the added gross profit.
Answer #1

(a) Fixed Cost = $300000

Selling Price = $75

Variable Cost = $50

Let the number of breakeven units be x

At breakeven point, Revenue = Costs

=> 75x = 300000 + 50x

=> x = 300000/(75-50) = 12,000 units

(b) Fixed Cost = $250000

Selling Price = $75

Variable Cost = $59

Let the number of breakeven units be x

At breakeven point, Revenue = Costs

=> 75x = 250000 + 59x

=> x = 250000/(75-59) = 15,625 units

(c) Fixed Cost = $250000

Selling Price = $75

Variable Cost = $59

Let the number of breakeven units be x

At breakeven point, Revenue = Costs

=> 75x + 150000 = 250000 + 59x

=> x = 100000/(75-59) = 6250 units